OpenStudy (anonymous):

verify. tanx + 1+secx ----- -------- = 2cscx 1+secx tanx

OpenStudy (anonymous):

\[\frac{\frac{b}{a}}{1+\frac{1}{a}}+\frac{1+\frac{1}{a}}{\frac{b}{a}}\]

OpenStudy (anonymous):